Operation Summary for June 14, 2010 - June 16, 2010 |
Monday June 14 The day shift began with Operations monitoring store 7891 (131.2E30) and the Recycler stash (162E10), with the Antiproton Source (Pbar) stacking (15.6E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and the Meson FTTB T-977 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations disabled stacking for 45 minutes to allow Pbar personnel to investigate a power supply ground fault. Operations also disabled the MiniBooNE beam for 40 minutes to allow FESS personnel to investigate air conditioning problems at MI-12. During the evening shift, Operations terminated store 7891 at 6:30 PM, put the TeV into HEP shot setup, and established store 7892 at 8 PM with an average initial luminosity (AIL) of 336.1E30. Operations disabled the T-977 beam at 6:07 PM. Tuesday June 15 The midnight shift began with Operations monitoring store 7892 (156.79E30) and the stash (145.9E10), with Pbar stacking (18.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported that a TeV RF station (TRF2) tripped off and caused a loss of 5.5E30 in luminosity. Operations established beam to T-977 at 4:13 AM. During the day shift, Operations terminated store 7892 at 9:30 AM. Operations established a TeV study store, 7893, with three bunches of protons and three bunches of antiprotons at 11:15 AM with an average initial luminosity of 11E30. Operations terminated store 7893 at 2:15 PM, put the TeV into shot setup, and established store 7894 at 3:42 PM with an AIL of 334E30. Operations turned off the MiniBooNE beam for air conditioning work. Operations halted stacking for Debuncher bend bus ground fault repairs. During the evening shift, Operations aborted store 7894 at 9 PM when a flashover in a transformer bus duct tripped off several TeV cryo system devices. There will be no T-977 beam until the transformer bus duct has been repaired. A thunderstorm held MI off with main bus power supply problems for three hours, which interrupted stacking and NuMI beam. Operations reestablished beam to MiniBooNE and NuMI by 11 PM and Pbar resumed stacking by the end of shift. The TeV recovery continued. Wednesday June 16 The midnight shift began with the TeV recovering from a power glitch, with Operations monitoring the stash (53.7E10) and the Pbar stack (12.9E10), with Pbar in standby due to ARF1 problems, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported that the TeV had fully recovered, but a TeV quadrupole power supply (B0Q6) had regulation problems; operators called in power supply personnel to make repairs. They had the supply repaired by 6:30 PM. Machine checkouts and tune-ups filled out the rest of the shift. The Plans for Wednesday The plans for today are to stack, stash, and store. The Run Coordinator will allow BRF6 work and MI studies. |
|
More Information Comments and Suggestions |
| last modified 06/16/2010 email Fermilab |
| Security, Privacy, Legal | |